Circular No. 5703

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

SUPERNOVA 1993D IN ESO 571-G3 R. H. McNaught, Bugaldie, N.S.W, reports his discovery of an apparent

supernova of mag ∼ 18 on an R plate taken on Feb. 1.7 UT by M. J. Drinkwater with the U.K. Schmidt Telescope. SN 1993D is located at α = 11^{h}26^{m}28^{s}.90, δ = −22^{o}12′34′′.0 (equinox 1950.0; uncertainty 0′′.5 in each coordinate), and it lies on the inner edge of the ring of ESO 571-G3 (= MCG –4-27-11), being offset 10′′.7 west and 6′′.8 south from the galaxy’s center. A nearby star of mag 17.5 has end figures 31^{s}.40, 03′′.2. No star appears in the position of the supernova on the ESO or SERC R surveys.

M. Della Valle and L. Pasquini, European Southern Observatory, report: “Inspection of a spectrogram (range 450-720 nm, resolution ∼ 2 nm) obtained on Feb. 3.3 UT with the ESO 1.5-m telescope (+ Boller & Chivens spectrograph) at La Silla shows this to be a type-Ia supernova caught ∼ 3 weeks after maximum.”

GRB 930131 T. L. Cline, S. Barthelmy, and D. Palmer, Goddard Space Flight Cen-

ter, NASA, on behalf of the Ulysses/Compton Interplanetary Network team for the localization of γ-ray burst sources (K. Hurley, M. Sommer, M. Boer, and M. Niel for Ulysses; G. Fishman, C. Kouveliotou, C. Meegan, W. Paciesas, and R. Wilson for CGRO/BATSE), communicate: “We report the source location of the event of Jan. 31.78972 UT. The source annulus defined by the observations from Ulysses and Compton Gamma Ray Observatory/BATSE has a radius of 22^{o}.253, from a center at α = 10^{h}43^{m}.2, δ = −14^{o}18′ (equinox J2000.0). This annulus intersects the CGRO/Comptel source region (published on IAUC 5702), and provides a ring segment 1′ thick passing through α = 12^{h}13^{m}52^{s}, δ = −11^{o}37′.3 and α = 12^{h}09^{m}08^{s}, δ = −7^{o}13′.0, reducing the IAUC 5702 source region by nearly two orders of magnitude. This is an exceptionally intense burst with an extremely rapid onset, and could be a good candidate for study at other wavelengths.”

PERIODIC COMET SCHAUMASSE (1992x) Corrigendum. On IAUC 5691, line 2, for 17.26, 12.7 read Dec. 17.26,

12.7

1993 February 3 Daniel W. E. Green